OpenStudy (haleyelizabeth2017):

Compare each function with the parent function. Without graphing, what are the vertex, axis of symmetry, and transformations of the parent function? y = –|8x + 4| + 2

OpenStudy (haleyelizabeth2017):

One sec and I will get the options.

OpenStudy (haleyelizabeth2017):

Possible answers: (1/2,2) ; x=-1/2 ; translated to the left 1/2 unit and up 2 units. (-1/2,-2) ; x=-1/2 ; translated to the left 1/2 unit, up 2 units, and reflected in the x-axis. (1/2,2) ; x=1/2 ; translated to the right 1/2 unit and up 2 units, and reflected in the y-axis. (1/2,-2) ; x=1/2 ; translated to the right 1/2 unit and up 2 units